# Fin-R1: A Large Language Model for Financial Reasoning through Reinforcement Learning

### v3

@@ Abstract
In recent years, general-purpose large language models (LLMs) such as GPT, Gemini, Claude, and DeepSeek have advanced at an unprecedented pace. Despite these achievements, their application to finance remains challenging, due to fragmented data sources, intransparent reasoning processes, and weak transferability to business applications. In response, we introduce Fin-R1, a reasoning LLM designed for financial scenarios. With a compact size of 7 billion parameters, Fin-R1 reduces deployment costs while addressing the aforementioned challenges. Its development follows a two-stage pipeline. First, we construct Fin-R1-Data, a high-quality financial dataset consisting of 60,091 chain-of-thought (CoT) samples, distilled and filtered from multiple authoritative benchmarks to ensure consistency and reliability. Second, we train Fin-R1 using Fin-R1-Data through supervised fine-tuning (SFT), followed by reinforcement learning (RL). This stage substantially improves the model’s ability to solve complex financial reasoning tasks, yielding outputs that are both accurate and interpretable. Despite its relatively small parameter scale, Fin-R1 achieves competitive empirical performance across established financial benchmarks and demonstrates practical utility in compliance checking and robo-advisory. Our code is publicly available at https://github.com/SUFE-AIFLM-Lab/Fin-R1, and has already attracted over 700 stars.
@@ 1 Introduction
Large language models (LLMs) are highly sophisticated neural networks that can understand and generate human language. In recent years, they have achieved groundbreaking progress in natural language processing, making an important step toward artificial general intelligence (AGI, touvron2023llama; team2024gemini; naveed2025comprehensive). At the core of LLMs lies a two-stage training paradigm. The first stage, pre-training (bai2023qwen; anil2023palm), leverages the transformer architecture (vaswani2017attention) to enable LLMs to learn linguistic patterns, world knowledge, and basic reasoning capacities from vast amounts of unlabeled text. However, pre-trained LLMs remain limited in their ability to handle complex reasoning tasks. This limitation has led to the second stage, post-training, where customized algorithms have been developed to enhance reasoning ability and improve output quality. These algorithms include supervised fine-tuning (SFT, ouyang2022training) to inject reasoning knowledge, reinforcement learning from human feedback (RLHF, christiano2017deep) to improve output quality, and reinforcement learning with verifiable rewards (RLVR, lambert2024tulu) to optimize the model’s reasoning capabilities through objectively verifiable signals. Although post-training is conducted on small-scale data and constitutes less than 1% of the total training computation (Lai_2025_Survey), it plays a pivotal role in enhancing reasoning performance, task alignment, and response reliability, enabling LLMs to handle various complex tasks at near-human levels.
Building on these post-training algorithms, a new generation of general-purpose, reasoning-oriented LLMs has emerged, including OpenAI’s o1 (chatgpt) and o1-like models such as QwQ (qwen) and Marco-o1 (zhao2024marco). These models explicitly incorporate an exploration–reflection–iteration (ERI) mechanism to enhance reasoning. In ERI, the model first explores multiple candidate reasoning traces, then verifies them through self-consistency or verifiable signals, and finally iterates to refine the reasoning process until convergence. This mechanism has led to significant improvements in chain-of-thought (CoT, wei2022chain) [^1] CoT refers to a sequence of intermediate reasoning steps that can be elicited either through carefully designed few-shot examples or through some simple magical prompts (e.g., “Let us think step by step”). reasoning on complex tasks, with performance often reaching the level of human experts. It is particularly useful in mathematical and logical problems that require a deep understanding of the problems themselves, step-by-step reasoning, and precise solutions. However, our empirical investigation reveals that such general-purpose, reasoning-oriented LLMs become limited when applied to the financial domain (see Section 5). These limitations stem primarily from three sources:
@@
    Financial data are highly fragmented and lack a unified structure (wang2024quantagent; guo2025fineval; li2024alphafin; dong2024fnspid), making the integration of knowledge extremely difficult. Specifically, information such as contractual terms, regulatory requirements, macroeconomic indicators, and market signals is often dispersed across heterogeneous sources, and can be inconsistent or even mutually contradictory. This not only increases the cost of data preprocessing but also explains why high-quality CoT data remain scarce in finance. As financial reasoning relies on the integration of economic, legal, and quantitative logic, the shortage of high-quality data limits the capability of existing models to reason coherently in finance.
@@
    Most existing LLMs still operate as “black boxes,” with reasoning processes that remain intransparent: we only observe the final outputs, but not the underlying reasoning paths (wang2023alpha; zhao2024explainability; tong2024ploutos). Such lack of transparency conflicts with the regulatory and compliance requirements of the financial sector where traceability and explainability are essential, restricting their deployment in practice.
@@
    In high-stakes financial applications, existing models often suffer from weak transferability and generalization, which makes their outputs unreliable (yu2024fincon; fatouros2024can; zhou2023universalner). Specifically, financial tasks involve evolving environments, which challenge the ability of existing models to generalize and adapt across scenarios. For instance, models trained solely via SFT rely heavily on memorizing previously seen examples rather than performing reasoning grounded in financial logic (zhang2025policy). This largely limits their effectiveness in applications such as credit assessment and risk pricing.
@@
  • In the first stage, we construct a reasoning dataset for finance, termed as Fin-R1-Data, which covers a wide range of financial scenarios encountered in practice. This stage addresses the first challenge of fragmented data sources and the scarcity of CoT data within financial domains, offering the basis for the subsequent SFT and RL. Specifically, Fin-R1-Data is a high-quality bilingual dataset containing over 60,000 entries, integrating data from diverse sources – including open-source datasets and proprietary examination problems – to ensure comprehensive coverage of professional expertise, business practices, and numerical reasoning. Its construction proceeds in two steps. During the data distillation step, we leverage DeepSeek-R1 (guo2025deepseek) to produce reasoning traces while standardizing answer formats. In the subsequent data filtering step, we employ Qwen2.5-72B-Instruct (yang2024qwen2) as an evaluator to investigate the logical consistency, coherence, and domain alignment of the data, screen out low-quality samples and retain high-quality data for reasoning. The left panel of Figure 1 depicts a high-quality CoT example after filtering.
  • In the second stage, we post-train LLMs by applying SFT and RL on Fin-R1-Data to obtain our Fin-R1 model. This stage substantially enhances the pre-trained model’s generalization and transferability across diverse business applications, addressing the third challenge. Specifically, SFT uses carefully collected CoT samples from Fin-R1-Data to enable the model to “think before answering” and to conduct integrated reasoning across legal, economic, and quantitative domains. The next step applies GRPO, a computationally efficient variant of proximal policy optimization (PPO, schulman2017proximal) which evaluates model’s outputs using group-relative advantage functions, to ensure both correctness and structural consistency in reasoning chains. Such a two-step post-training approach – first SFT, then GRPO – enables a 7B-parameter model to deliver reliable and trustworthy outputs in financial applications. Notably, despite being 100 times smaller than leading frontier reasoning models, Fin-R1 achieved an average score of 75.2 on financial reasoning benchmarks, ranking second overall and delivering the nearly best performance on mainstream financial benchmarks (e.g., FinQA, chen-etal-2021-finqaa). Compared with models of the same 7B scale, Fin-R1 outperformed the existing state-of-the-art models by more than 17 points.
@@ 2 Related Work
Pre-training. Pre-training enables LLMs to learn linguistic patterns and fundamental knowledge from massive text corpora (devlin2019bert; floridi2020gpt; xue2020mt5). Through auto-regressive next-token prediction on trillions of internet tokens, these models absorbed knowledge across diverse domains such as programming (kocetkov2022stack), mathematics (drori2022neural), and other professional or scientific fields (lo2019s2orc). However, LLM pre-training faces several practical challenges. First, it is highly resource-intensive: for example, training PaLM with 540 billion parameters requires 6,144 TPU v4 chips over several weeks (chowdhery2023palm). Second, as high-quality internet text data show signs of exhaustion (villalobos2022will), the scaling law  (kaplan2020scaling) that links model size, data volume, and performance appears to approach its limit. Although large-scale synthetic corpora can be potentially employed, their effectiveness has not yet been fully validated (shumailov2024ai). Third, pre-trained models often struggle to generalize to domain-specific reasoning tasks. This hinders their abilities to understand complex financial logic, such as derivative pricing or risk-hedging strategies (zoph2020rethinking; abnar2021exploring).
@@
Post-training. There is a growing consensus that additional algorithmic paradigms beyond pre-training are required to move closer to AGI. Recent reasoning-oriented models, such as OpenAI’s o1 and DeepSeek-R1, position post-training as such a paradigm (see Kumar_2025_LLM, for an overview). Post-training aims to mitigate pre-trained models’ limitations by aligning their outputs with human values, while reducing biases and inaccuracies (bai2022constitutional). It typically includes fine-tuning for task-specific adaptation (trung2024reft), RL for encouraging the model to explore better outputs guided by feedback signals (Zhang_2025_Survey), and test-time scaling for boosting model performance without retraining the model (TTS, hu2022lora). Among these methods, RL-based post-training has repeatedly demonstrated that well-defined reward signals can drive AI agents to achieve superhuman performance on complex tasks (ouyang2022training).
@@
Among RL algorithms, PPO, a computationally efficient approximation to the trust region policy optimization algorithm (schulman2015trust), has been widely adopted for post-training LLMs. By limiting the divergence between the old and new policies, PPO effectively mitigates the issue of policy collapse commonly encountered in traditional policy gradient algorithms. However, PPO suffers from several limitations, including training instability, high sensitivity to hyperparameter tuning, and substantial computational cost arising from the need to learn a value function for the language model (engstrom2020implementation; zheng2023secrets; xu2024dpo).
@@
To address these challenges, rafailov2023direct proposed a direct preference optimization (DPO), which expresses the reward model in closed-form using the optimal policy, and transforms the complex RL problem into a more tractable classification task. DPO eliminates the need for value function training, substantially lowers the computational cost, and reduces the number of hyperparameters. Nevertheless, it remains highly dependent on the reference policy and suffers from limited generalization to out-of-distribution samples (xu2024contrastive; feng2024towards; xu2025doubly; ye2025robust).
@@
Building on PPO, shao2024deepseekmath achieved a major breakthrough in LLM reasoning by introducing the GRPO algorithm. Unlike traditional PPO, which primarily relies on a single-stream reward to guide model optimization, GRPO generates multiple candidate outputs and computes advantages based on their relative performance within the group, thus eliminating the need for learning a value network and considerably improving the computational efficiency. Compared to both PPO and DPO, GRPO demonstrates clear advantages in reasoning-intensive tasks such as mathematical problem solving (guo2025deepseek). See Zhang_2025_Survey for a comprehensive discussion of RL algorithms for LLMs.
@@
Financial LLMs. Pre-trained and post-trained models discussed above are primarily tailored for general-purpose tasks. Recently, several financial LLMs have been developed, such as BloombergGPT (wu2023bloomberggpt), DISC-FinLLM (chen2023disc), PIXIU (xie2023pixiu), and XuanYuan (zhang2023xuanyuan). However, these models are primarily designed for non-reasoning tasks. In contrast, our proposed model, Fin-R1, can reason, which allows it to integrate fragmented financial knowledge across business applications to generate practically useful outputs. Meanwhile, models such as XuanYuan-FinX1-Preview (XuanYuan2024Duxiaoman) and Fino1 (qian2025fino1) adopt o1-like reasoning paradigm to strengthen numerical computation and logical analysis in financial applications. In parallel, FinAgent (zhang2024multimodal) enhances trading decision making via multimodal tool integration. These efforts collectively push financial LLMs from mere text understanding toward explicit reasoning and decision making. Nevertheless, there remains substantial room to improve reasoning accuracy and transparency, and to enhance model transferability across financial scenarios (huang2024open; xie2024finben).
@@ 3.1 Data Overview
Financial professional knowledge. This category of data is designed to handle complex, specialized application scenarios, covering questions that can only be answered by humans with sufficient professional financial expertise. It mainly contains three types of data: (i) explanations of financial terminology, providing precise interpretations of specialized terms and concepts; (ii) financial calculations, involving specialized computational algorithms and models; and (iii) questions from finance-related postgraduate entrance examinations. The last type of data is scarce. To address this, we constructed a financial postgraduate entrance exam (FinPEE) dataset using original examination questions from Shanghai University of Finance and Economics. This dataset was constructed in two steps: (a) text extraction, where PDF-based source materials were converted into Markdown format using Mineru (wang2024mineru), and (b) expert review, in which structured question–answer pairs were verified by financial experts to ensure accuracy.
@@ 3.2 Data Processing
Specifically, we first review the responses generated by DeepSeek-R1-671B during data distillation. For objective questions, we retain only those responses whose answers exactly match the correct ones. For subjective questions, we adopt the LLM-as-Judges paradigm – using a well-tuned LLM as an automated evaluator – to access the correctness of each answer and discard any incorrect ones. We further conduct a comparative study to compare GPT-4o (openai2024gpt4o) against Qwen2.5-72B-Instruct (yang2024qwen2) as judges, and evaluate different prompting strategies for enabling these LLMs to serve effectively in this role. Our experimental results show that Qwen2.5-72B-Instruct achieves an accuracy rate of 99.6% and outperforms GPT-4o. We thus select Qwen2.5-72B-Instruct as the judge model and employ the optimal prompt that achieves the highest accuracy. Further details on the experimental setup and results are provided in Section C.1 of the Supplementary Material.
We next focus on the filtered dataset that contains correct answers to the questions and evaluate their reasoning paths. Following xie2024finnlpa, we similarly adopt the LLM-as-Judges paradigm and prompt the LLM to evaluate the reasoning paths according to criteria from the following seven dimensions: internal consistency, term overlap rate, number of reasoning steps, logical coherence, content diversity, task-domain relevance, and alignment with task instructions. To evaluate the performance of LLMs as judges, we conducted additional experiments to compare human annotations against LLM evaluations. The results, detailed in Section C.2 of the Supplementary Material, demonstrate that Qwen2.5-72B-Instruct closely aligns with human judgments, exhibiting only minor discrepancies, whereas GPT-4o shows larger deviations. Based on these findings, we select Qwen2.5-72B-Instruct to assess the quality of the reasoning paths, and retain only the high-quality ones for the subsequent SFT. In the yellow boxes of Figure 4, we present examples of both a high-quality and a low-quality reasoning trajectory to illustrate their differences. The sentence beginning with “Thinking:” is generated by Qwen2.5-72B-Instruct, which analyzes the rationality of each reasoning trajectory and assigns a corresponding quality score.

### v4

@@ Abstract
In recent years, general-purpose large language models (LLMs) such as GPT, Gemini, Claude, and DeepSeek have advanced at an unprecedented pace. Despite these achievements, their application to finance remains challenging, due to fragmented data sources, reasoning processes, and weak transferability to business applications. In response, we introduce Fin-R1, a reasoning LLM designed for financial scenarios. With a compact size of 7 billion parameters, Fin-R1 reduces deployment costs while addressing the aforementioned challenges. Its development follows a two-stage pipeline. First, we construct Fin-R1-Data, a high-quality financial dataset consisting of 60,091 chain-of-thought (CoT) samples, distilled and filtered from multiple authoritative benchmarks to ensure consistency and reliability. Second, we train Fin-R1 using Fin-R1-Data through supervised fine-tuning (SFT), followed by reinforcement learning (RL). This stage substantially improves the model’s ability to solve complex financial reasoning tasks, yielding outputs that are both accurate and interpretable. Despite its relatively small parameter scale, Fin-R1 achieves competitive empirical performance across established financial benchmarks and demonstrates practical utility in compliance checking and robo-advisory. Our code is publicly available at https://github.com/SUFE-AIFLM-Lab/Fin-R1, and has already attracted over 700 stars.
@@ 1 Introduction
Large language models (LLMs) are highly sophisticated neural networks that can understand and generate human language. In recent years, they have achieved groundbreaking progress in natural language processing, making an important step toward artificial general intelligence (AGI, Touvron et al., 2023; Google, 2024; Naveed et al., 2025). At the core of LLMs lies a two-stage training paradigm. The first stage, pre-training (Bai et al., 2023; Google, 2023), leverages the transformer architecture (Vaswani et al., 2017) to enable LLMs to learn linguistic patterns, world knowledge, and basic reasoning capacities from vast amounts of unlabeled text. However, pre-trained LLMs remain limited in their ability to handle complex reasoning tasks. This limitation has led to the second stage, post-training, where customized algorithms have been developed to enhance reasoning ability and improve output quality. These algorithms include supervised fine-tuning (SFT, Ouyang et al., 2022) to inject reasoning knowledge, reinforcement learning from human feedback (RLHF, Christiano et al., 2017) to improve output quality, and reinforcement learning with verifiable rewards (RLVR, Lambert et al., 2024) to optimize the model’s reasoning capabilities through objectively verifiable signals. Although post-training is conducted on small-scale data and constitutes less than 1% of the total training computation (Lai et al., 2025), it plays a pivotal role in enhancing reasoning performance, task alignment, and response reliability, enabling LLMs to handle various complex tasks at near-human levels.
Building on these post-training algorithms, a new generation of general-purpose, reasoning-oriented LLMs has emerged, including OpenAI’s o1 (OpenAI, 2024b) and o1-like models such as QwQ (Team, 2024c) and Marco-o1 (Zhao et al., 2024b). These models explicitly incorporate an exploration–reflection–iteration (ERI) mechanism to enhance reasoning. In ERI, the model first explores multiple candidate reasoning traces, then verifies them through self-consistency or verifiable signals, and finally iterates to refine the reasoning process until convergence. This mechanism has led to significant improvements in chain-of-thought (CoT, Wei et al., 2022) [^1] CoT refers to a sequence of intermediate reasoning steps that can be elicited either through carefully designed few-shot examples or through some simple magical prompts (e.g., “Let us think step by step”). reasoning on complex tasks, with performance often reaching the level of human experts. It is particularly useful in mathematical and logical problems that require a deep understanding of the problems themselves, step-by-step reasoning, and precise solutions. However, our empirical investigation reveals that such general-purpose, reasoning-oriented LLMs become limited when applied to the financial domain (see Section 5). These limitations stem primarily from three sources:
@@
    Financial data are highly fragmented and lack a unified structure (Wang et al., 2024b; Guo et al., 2025b; Li et al., 2024; Dong et al., 2024), making the integration of knowledge extremely difficult. Specifically, information such as contractual terms, regulatory requirements, macroeconomic indicators, and market signals is often dispersed across heterogeneous sources, and can be inconsistent or even mutually contradictory. This not only increases the cost of data preprocessing but also explains why high-quality CoT data remain scarce in finance. As financial reasoning relies on the integration of economic, legal, and quantitative logic, the shortage of high-quality data limits the capability of existing models to reason coherently in finance.
@@
    Most existing LLMs still operate as “black boxes,” with reasoning processes that remain intransparent: we only observe the final outputs, but not the underlying reasoning paths (Wang et al., 2023; Zhao et al., 2024a; Tong et al., 2024). Such lack of transparency conflicts with the regulatory and compliance requirements of the financial sector where traceability and explainability are essential, restricting their deployment in practice.
@@
    In high-stakes financial applications, existing models often suffer from weak transferability and generalization, which makes their outputs unreliable (Yu et al., 2024; Fatouros et al., 2024; Zhou et al., 2024). Specifically, financial tasks involve evolving environments, which challenge the ability of existing models to generalize and adapt across scenarios. For instance, models trained solely via SFT rely heavily on memorizing previously seen examples rather than performing reasoning grounded in financial logic (Zhang et al., 2025b). This largely limits their effectiveness in applications such as credit assessment and risk pricing.
@@
  • In the first stage, we construct a reasoning dataset for finance, termed as Fin-R1-Data, which covers a wide range of financial scenarios encountered in practice. This stage addresses the first challenge of fragmented data sources and the scarcity of CoT data within financial domains, offering the basis for the subsequent SFT and RL. Specifically, Fin-R1-Data is a high-quality bilingual dataset containing over 60,000 entries, integrating data from diverse sources – including open-source datasets and proprietary examination problems – to ensure comprehensive coverage of professional expertise, business practices, and numerical reasoning. Its construction proceeds in two steps. During the data distillation step, we leverage DeepSeek-R1 (Guo et al., 2025a) to produce reasoning traces while standardizing answer formats. In the subsequent data filtering step, we employ Qwen2.5-72B-Instruct (Yang et al., 2024) as an evaluator to investigate the logical consistency, coherence, and domain alignment of the data, screen out low-quality samples and retain high-quality data for reasoning. The left panel of Figure 1 depicts a high-quality CoT example after filtering.
  • In the second stage, we post-train LLMs by applying SFT and RL on Fin-R1-Data to obtain our Fin-R1 model. This stage substantially enhances the pre-trained model’s generalization and transferability across diverse business applications, addressing the third challenge. Specifically, SFT uses carefully collected CoT samples from Fin-R1-Data to enable the model to “think before answering” and to conduct integrated reasoning across legal, economic, and quantitative domains. The next step applies GRPO, a computationally efficient variant of proximal policy optimization (PPO, Schulman et al., 2017) which evaluates model’s outputs using group-relative advantage functions, to ensure both correctness and structural consistency in reasoning chains. Such a two-step post-training approach – first SFT, then GRPO – enables a 7B-parameter model to deliver reliable and trustworthy outputs in financial applications. Notably, despite being 100 times smaller than leading frontier reasoning models, Fin-R1 achieved an average score of 75.2 on financial reasoning benchmarks, ranking second overall and delivering the nearly best performance on mainstream financial benchmarks (e.g., FinQA, Chen et al., 2021). Compared with models of the same 7B scale, Fin-R1 outperformed the existing state-of-the-art models by more than 17 points.
@@ 2 Related Work
Pre-training. Pre-training enables LLMs to learn linguistic patterns and fundamental knowledge from massive text corpora (Devlin et al., 2019; Floridi and Chiriatti, 2020; Xue et al., 2021). Through auto-regressive next-token prediction on trillions of internet tokens, these models absorbed knowledge across diverse domains such as programming (Kocetkov et al., 2022), mathematics (Drori et al., 2022), and other professional or scientific fields (Lo et al., 2020). However, LLM pre-training faces several practical challenges. First, it is highly resource-intensive: for example, training PaLM with 540 billion parameters requires 6,144 TPU v4 chips over several weeks (Chowdhery et al., 2023). Second, as high-quality internet text data show signs of exhaustion (Villalobos et al., 2024), the scaling law  (Kaplan et al., 2020) that links model size, data volume, and performance appears to approach its limit. Although large-scale synthetic corpora can be potentially employed, their effectiveness has not yet been fully validated (Shumailov et al., 2024). Third, pre-trained models often struggle to generalize to domain-specific reasoning tasks. This hinders their abilities to understand complex financial logic, such as derivative pricing or risk-hedging strategies (Zoph et al., 2020; Abnar et al., 2021).
@@
Post-training. There is a growing consensus that additional algorithmic paradigms beyond pre-training are required to move closer to AGI. Recent reasoning-oriented models, such as OpenAI’s o1 and DeepSeek-R1, position post-training as such a paradigm (see Kumar et al., 2025, for an overview). Post-training aims to mitigate pre-trained models’ limitations by aligning their outputs with human values, while reducing biases and inaccuracies (Bai et al., 2022). It typically includes fine-tuning for task-specific adaptation (Trung et al., 2024), RL for encouraging the model to explore better outputs guided by feedback signals (Zhang et al., 2025a), and test-time scaling for boosting model performance without retraining the model (TTS, Hu et al., 2022). Among these methods, RL-based post-training has repeatedly demonstrated that well-defined reward signals can drive AI agents to achieve superhuman performance on complex tasks (Ouyang et al., 2022).
@@
Among RL algorithms, PPO, a computationally efficient approximation to the trust region policy optimization algorithm (Schulman et al., 2015), has been widely adopted for post-training LLMs. By limiting the divergence between the old and new policies, PPO effectively mitigates the issue of policy collapse commonly encountered in traditional policy gradient algorithms. However, PPO suffers from several limitations, including training instability, high sensitivity to hyperparameter tuning, and substantial computational cost arising from the need to learn a value function for the language model (Engstrom et al., 2020; Zheng et al., 2023; Xu et al., 2024b).
@@
To address these challenges, Rafailov et al. (2023) proposed a direct preference optimization (DPO), which expresses the reward model in closed-form using the optimal policy, and transforms the complex RL problem into a more tractable classification task. DPO eliminates the need for value function training, substantially lowers the computational cost, and reduces the number of hyperparameters. Nevertheless, it remains highly dependent on the reference policy and suffers from limited generalization to out-of-distribution samples (Xu et al., 2024a; Feng et al., 2024; Xu et al., 2025; Ye et al., 2025).
@@
Building on PPO, Shao et al. (2024) achieved a major breakthrough in LLM reasoning by introducing the GRPO algorithm. Unlike traditional PPO, which primarily relies on a single-stream reward to guide model optimization, GRPO generates multiple candidate outputs and computes advantages based on their relative performance within the group, thus eliminating the need for learning a value network and considerably improving the computational efficiency. Compared to both PPO and DPO, GRPO demonstrates clear advantages in reasoning-intensive tasks such as mathematical problem solving (Guo et al., 2025a). See Zhang et al. (2025a) for a comprehensive discussion of RL algorithms for LLMs.
@@
Financial LLMs. Pre-trained and post-trained models discussed above are primarily tailored for general-purpose tasks. Recently, several financial LLMs have been developed, such as BloombergGPT (Wu et al., 2023), DISC-FinLLM (Chen et al., 2023), PIXIU (Xie et al., 2023), and XuanYuan (Zhang and Yang, 2023). However, these models are primarily designed for non-reasoning tasks. In contrast, our proposed model, Fin-R1, can reason, which allows it to integrate fragmented financial knowledge across business applications to generate practically useful outputs. Meanwhile, models such as XuanYuan-FinX1-Preview (Team, 2024a) and Fino1 (Qian et al., 2025) adopt o1-like reasoning paradigm to strengthen numerical computation and logical analysis in financial applications. In parallel, FinAgent (Zhang et al., 2024) enhances trading decision making via multimodal tool integration. These efforts collectively push financial LLMs from mere text understanding toward explicit reasoning and decision making. Nevertheless, there remains substantial room to improve reasoning accuracy and transparency, and to enhance model transferability across financial scenarios (Huang et al., 2024; Xie et al., 2024a).
@@ 3.1 Data Overview
Financial professional knowledge. This category of data is designed to handle complex, specialized application scenarios, covering questions that can only be answered by humans with sufficient professional financial expertise. It mainly contains three types of data: (i) explanations of financial terminology, providing precise interpretations of specialized terms and concepts; (ii) financial calculations, involving specialized computational algorithms and models; and (iii) questions from finance-related postgraduate entrance examinations. The last type of data is scarce. To address this, we constructed a financial postgraduate entrance exam (FinPEE) dataset using original examination questions from Shanghai University of Finance and Economics. This dataset was constructed in two steps: (a) text extraction, where PDF-based source materials were converted into Markdown format using Mineru (Wang et al., 2024a), and (b) expert review, in which structured question–answer pairs were verified by financial experts to ensure accuracy.
@@ 3.2 Data Processing
Specifically, we first review the responses generated by DeepSeek-R1-671B during data distillation. For objective questions, we retain only those responses whose answers exactly match the correct ones. For subjective questions, we adopt the LLM-as-Judges paradigm – using a well-tuned LLM as an automated evaluator – to access the correctness of each answer and discard any incorrect ones. We further conduct a comparative study to compare GPT-4o (OpenAI, 2024a) against Qwen2.5-72B-Instruct (Yang et al., 2024) as judges, and evaluate different prompting strategies for enabling these LLMs to serve effectively in this role. Our experimental results show that Qwen2.5-72B-Instruct achieves an accuracy rate of 99.6% and outperforms GPT-4o. We thus select Qwen2.5-72B-Instruct as the judge model and employ the optimal prompt that achieves the highest accuracy. Further details on the experimental setup and results are provided in Section C.1 of the Supplementary Material.
We next focus on the filtered dataset that contains correct answers to the questions and evaluate their reasoning paths. Following Xie et al. (2024b), we similarly adopt the LLM-as-Judges paradigm and prompt the LLM to evaluate the reasoning paths according to criteria from the following seven dimensions: internal consistency, term overlap rate, number of reasoning steps, logical coherence, content diversity, task-domain relevance, and alignment with task instructions. To evaluate the performance of LLMs as judges, we conducted additional experiments to compare human annotations against LLM evaluations. The results, detailed in Section C.2 of the Supplementary Material, demonstrate that Qwen2.5-72B-Instruct closely aligns with human judgments, exhibiting only minor discrepancies, whereas GPT-4o shows larger deviations. Based on these findings, we select Qwen2.5-72B-Instruct to assess the quality of the reasoning paths, and retain only the high-quality ones for the subsequent SFT. In the yellow boxes of Figure 4, we present examples of both a high-quality and a low-quality reasoning trajectory to illustrate their differences. The sentence beginning with “Thinking:” is generated by Qwen2.5-72B-Instruct, which analyzes the rationality of each reasoning trajectory and assigns a corresponding quality score.